OpenStudy (anonymous):

help solving systems of equations with three variables 1. 2a+c=-10, b-c=15, a-2b+c=-5

OpenStudy (anonymous):

use method of substitution

OpenStudy (anonymous):

ok what do you do after 2a+b=5 and a-3b=10

OpenStudy (anonymous):

so a= (5-b)/2 put it in a-3b=10 in place of a you will get b

OpenStudy (anonymous):

ok